• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name R. Michaels

    Log entry time 00:57:02 on October 9,2009

    Entry number 293918

    keyword=Compton update: no Compton tonight

    No Compton tonight !
    Megan and I went over to MCC to get the story straight.
    The injector/optics experts were here during swing; they 
    tried but did not yet succeed to reduce the background 
    to an acceptable level.  The HV trips every 2 min at 1450 V.  
    MCC says it's not possible to improve the tune tonight.  
    Will continue tomorrow at ~9 am.  Alex is in town (just
    showed up) and will overlap with them during the day.
    Edet is in garage.  HV off.  We will run regular HAPPEX DAQ
    production tonight.